A store advertises an air conditioner on sale for $280. If this is marked down 20%, what was the original price.
Again I tried to use a proportion or multiply 20% x 280 and I do not get the answer in my book. Please help me.

Found 2 solutions by Earlsdon, glabow:
Answer by Earlsdon(6103)

Sale price = original price - discount. Let $x = original price. 20% = 0.2
$280 = $x - $0.2x
$280 = $x(1-0.2)
$280 = $0.8x Divide both sides by 0.8
$350 = $x
Original price was $350.
Check:
$350 - 0.2($350)
$350 - $70 = $280

Answer by glabow(165)

Let x be the original price.
x - .2x = 280
.8x = 280
x = 280 / .8 = 280 x 10/8
x = 2800/8 = 350
Checking:
350 - .2 x 350 = 350 - 70 = 280